OpenStudy (anonymous):

Esther Flowers lives in a state which charges a car registration fee of $15.50 for the first 2,000 pounds and $.75 for every 100 pounds or fraction of 100 pounds above 2,000. Her car weighs 2,850 pounds. Find the registration fee.

OpenStudy (anonymous):

i think this right since 15.00 is for 2,00 pounds u dont have to find dat just the rest so 2,000-2850=850 .75x850=637.50 and add the 15,000 to this ans 637.50+15.50=653

OpenStudy (anonymous):

Well I would have thought that this process is right but I already know the answer is $22.25

OpenStudy (anonymous):

Does anyone see what I mean? :-)

OpenStudy (anonymous):

@DLS @amistre64

OpenStudy (anonymous):

is Arandas way right and am I missing it or is my answer right? :0

OpenStudy (amistre64):

Esther Flowers lives in a state which charges a car registration fee of $15.50 for the first 2,000 pounds $.75 for every 100 pounds or fraction of 100 pounds above 2,000. Her car weighs 2,850 pounds. Find the registration fee. 2850 = 2000 + 850 15.50 + 850 850 = 8.5 hundred pouds 15.50 + 8.5(.75) = reg fee

OpenStudy (anonymous):

Thank you! :-)

OpenStudy (amistre64):

youre welcome